# Additivity of Crossing Number via Restricted Reidemeister Moves

### v1

@@ 1 Introduction
Given a knot diagram $K$, denote by $\chi(K)$ the number of crossings in that diagram. Denote by $\mathcal{R}$ the set of classical Reidemeister moves and Reidemeister equivalence between diagrams as $K\sim_{\mathcal{R}}K^{\prime}$. Then $c(K)$ is defined as the minimim of $\{\chi(K^{\prime})\,:\,K^{\prime}\sim_{\mathcal{R}}K\}$. Now the question of additivity can be reformulated as follows. Is it possible that $K^{\prime}\sim_{\mathcal{R}}K_{0}\mathbin{\#}K_{1}$ and $\chi(K^{\prime})<c(K_{0})+c(K_{1})$? We will define a set of restricted Reidemeister moves $\mathcal{R}^{\prime}$ (in fact several versions) and show that if $K^{\prime}\sim_{\mathcal{R}^{\prime}}K$, then $c(K^{\prime})\geqslant c(K_{0})+c(K_{1})$.
@@ 2.1 Basics
The composition of functions is denoted $f\circ g$ but sometimes $\circ$” is ignored and it is denoted fgfg. If AA is a set, denote by $fA=f[A]=f(A)=|f\!\restriction\!A|=\{f(x)\,:\,x\in A\}$ the image of AA (usually $f[A]$ is preferred). By $\operatorname{id}_{A}$ we denote the identity map $A\to A$. The identity map fixes AA pointwise. Sometimes we say that a function $f$ fixes a set AA pointwise to mean that $f\!\restriction\!A=\operatorname{id}_{A}$.
By $\overset{\circ}{U}{}$, $\partial U$, and $\bar{U}$ denote the interior, closure, and the boundary of a set $U\subset\mathbb{R}^{n}$ respectively. By $\operatorname{Card}(A)$ denote the cardinality of the set AA. If $\gamma_{1}$ and $\gamma_{2}$ are mappings from II into some space $X$ such that $\gamma_{1}(1)=\gamma_{2}(0)$, we denote by $\gamma_{1}{}^{\frown}\gamma_{2}$ the map $\gamma\colon I\to\mathbb{R}^{2}$ given by $\gamma(t)=\gamma_{1}(2t)$ if $t<1/2$, and $\gamma(t)=\gamma_{2}(2t-1)$ otherwise. By $B(x,r)$ and $\overset{\circ}{B}{}(x,r)$ denote the closed and open balls respectively centered at $x$ with radius $r$.
@@ 2.2 Genericity
A polyhedron in $\mathbb{R}^{n}$ is a set which is the realization of a finite simplicial complex in $\mathbb{R}^{n}$ compatible with the linear structure of $\mathbb{R}^{n}$. A map $f\colon X\to Y$ between polyhedra $X$ and $Y$ is an immersion if for all $x\in X$ there is a polyhedral neighborhood $U$ of $x$ (i.e. $U$ is a polyhedron and $x\in\overset{\circ}{U}{}$) such that $f\!\restriction\!U$ is an embedding. A point $x\in|f|$ is a multiplicity point if $\operatorname{Card}(f^{-1}(x))>1$. It is a double point if $\operatorname{Card}(f^{-1}(x))=2$. Suppose $T\colon D\to\mathbb{R}^{2}$ is an immersion where $D$ is a 11-manifold with boundary. Suppose that $x\in\mathbb{R}^{2}$ is a double point, Then $x$ is transversal if there is a compact polyhedral neighborhood $U$ of $x$, a homeomorphism $h\colon U\to[-1,1]^{2}$, and disjoint closed connected sets $C_{1},C_{2}\subset D$ each of which contains one of the points of $T^{-1}(x)$ in their interior such that $|h\circ(T\!\restriction\!C_{1})|=[-1,1]\times\{0\}$ and $|h\circ(T\!\restriction\!C_{2})|=\{0\}\times[-1,1]$.
Definition $2.$
Let $D$ a 11-manifold with boundary. We say that an immersion $g$ of $D$ into $\mathbb{R}^{2}$ is generic if there are at most finitely many multiplicity points, all of them are transversal double points and none of them is at $g[\partial D]$. These double points are called crossings and the set of crossings is denoted
@@ 3.3 Composition of Links
A version of Theorem 3.1 also holds for links in the following way. A link diagram is a triple $L=(L,M,\ell)$ where $L\colon D\to\square$ is a generic immersion, $M\in\mathbb{N}$, $D=I\cup(M\times S^{1})$ is the disjoint union of II with $M$ many circles, $L\!\restriction\!I=(L\!\restriction\!I,\ell\!\restriction\!\operatorname{Cross}(L\!\restriction\!I))$ is a knot diagram, and for all $t\in M\times S^{1}$, $L(t)\in\overset{\circ}{\square}{}$. It is centered, if $L\!\restriction\!I$ is centered. It is composite, if $L\cap H=\{o\}$. Note that it follows that $(L\!\restriction\!I)\cap H=\{o\}$ is composite (recall the definition of $H$ in the beginning of Section 2.6). A partitioned link is a tuple $(L,M^{+},M^{-},\ell)$ such that $(L,M^{+}\cup M^{-},\ell)$ is a link and $M=M^{+}\sqcup M^{-}$. For a composite $L$, there is a unique partition $M=M^{+}\sqcup M^{-}$ such that
@@ 4.4 Applying a Homeomorphism to a Tangle
Suppose $h\colon\square\to\square$ is a homeomorphism and $(T,N,M,\ell)$ a t-diagram. Then we denote by hThT the t-diagram $(h\circ T,N,M,\ell\circ h^{-1})$. Important special cases are the applications of the moves $R_{0}$ and $R^{*}_{0}$, (Section 2.5).
@@ 4.5 Reparametrization
Sometimes t-diagrams need to be reparametrized. An example of this was the restriction of a t-diagram to a rectangle $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$, Definition $4.$ Another example is smoothing. A local smoothing of a t-diagram is a local change at a crossing as depicted in Figure 6. If t-diagrams $(T,N,M,\ell)$ and $(T^{\prime},N^{\prime},M^{\prime},\ell^{\prime})$ are the same t-diagram before and after smoothing, the number of strands and loops may or may not change, i.e. we might have $N\neq N^{\prime}$ and/or $M\neq M^{\prime}$, or they might not change etc. We will not give the technical details of the reparametrization and assume that the reader understands how it can be done in each specific case. A smoothing operation may connect two strands which were originally oriented in opposite directions. Then one of them changes orientation after merging with the other strand. This is not a problem, as in this context we do not care about the orientation of strands and loops in t-diagrams.
Similarly for curves. We introduced the notations $\gamma_{1}{}^{\frown}\gamma_{2}$ and $f\cup g$ in the beginning of this section, but sometimes the domain of the curves may not be exactly equal to II, or they may be oriented in the “wrong” direction, i.e. $\gamma_{1}(0)=\gamma_{2}(0)$ and we want to connect them at their beginning points. It should always be clear from the context how to (re-)parametrize the curves in a proper manner.
@@ 5.1 The Minimal Ray Intersection Number
A sequence of $R_{0}$-moves is equivalent to one $R_{0}$-move, because the composition of witnessing homeomorphisms is a witnessing homeomorphism. So suppose that $\widehat{K}$ has been obtained from $K$ using such a homeomorphism $h$. Recall that by the definition of the $R_{0}$-move, $h$ keeps $o$ fixed, $J$ is a ray if and only if hJhJ is a ray and $\operatorname{\mathbf{i}}(K^{\pm},J)=\operatorname{\mathbf{i}}(hK^{\pm},hJ)$. So the statement follows from the obvious equality $hK^{\pm}=\widehat{K}^{\pm}$. ∎
@@ 6.1 Multigraphs, Matrices and Zigzag Paths
Denote by $\mathbb{N}$ natural numbers including zero and by $\mathbb{N}_{+}=\mathbb{N}\setminus\{0\}$ natural numbers excluding zero. For $n,m\in\mathbb{N}$ denote by $\llbracket n,m\rrbracket$ the discrete set $\{n,n+1,\dots,m-1,m\}$. By convention, if $m<n$, then $\llbracket n,m\rrbracket$ is the empty set and if $m=n$, then it is the singleton $\{n\}=\{m\}$. Recall that $\operatorname{Card}(A)$ the cardinality of the set AA, e.g. for $n\leqslant m$ we have
@@
The ijij-th element of an $(n\times n)$-matrix $e$ is expressed as $e(i,j)$ where $i,j\in\llbracket 1,n\rrbracket$. Let $\mathcal{M}^{s}_{n}(\mathbb{N})$ be the set of symmetric $\mathbb{N}$-valued $(n\times n)$-matrices,
@@
In the multigraph interpretation, $s_{i}(e)$ is the total number of out-going connections from ii.
@@
Clearly $x_{q}>x_{q+1}$. By (6.5), ($enumi^{m}_{3}$), and the symmetry of $x_{q}$ we have $x_{q+1}\in\mathcal{M}^{s}_{n}(\mathbb{N})$. It remains to verify that $x_{q+1}$ satisfies $(*^{m^{\prime}})$. The condition $(*^{m^{\prime}}_{3})$ for $x_{q+1}$ follows trivially from the assumption that ($enumi^{m}_{3}$) holds for $x_{q}$. So let us prove $(*^{m^{\prime}}_{1})$ and $(*^{m^{\prime}}_{2})$ for $x_{q+1}$. Suppose first that $f(m)\neq m^{\prime}$. Then the definition of $x_{q+1}$ means that the number 11 has been subtracted from exactly two rows in $x_{q}$ to obtain $x_{q+1}$. These rows are $f(m)$ and $m^{\prime}$. This means that
@@
We have proved $(*^{m^{\prime}})$ for $x_{q+1}$ in the case $m^{\prime}\neq f(m)$. Let us now assume that $f(m)=m^{\prime}$. Then the definition of $x_{q+1}$ means that the number 22 has been subtracted frmo exactly one row of $x_{q}$ to obtain $x_{q+1}$ that row being $m^{\prime}=f(m)$. This means:
@@
Given two matrices $a$ and $b$, denote by abab their Hadamard product, i.e. the matrix such that $(ab)(i,j)=a(i,j)b(i,j)$. Let us also introduce temporarily the following shorthand notation for a matrix $a$ and a set AA: ${(A)a}=\sum_{(i,j)\in A}a(i,j)$. Then the statement (6.24) becomes
@@ 6.2 Minimal Splitting Curves
Definition $6.$
Recall Definition 4.6 of $L_{1}(T),\dots,L_{m}(T)$. Given a t-diagram $T$ and a splitting curve $J$, if the starting point of $J$ is in some $L_{i}(T)$ and the ending point is in some $L_{j}(T)$, then we call $J$ an ijij-splitting curve (of $T$).
@@
Given a t-diagram $T=(T,N,M,\ell)$ denote by $\operatorname{MIN}_{ij}(T)$ the minimum number of intersections that an ijij-splitting curve can have with $T$, formally
@@
If $m=0$ then let $\beta=\varnothing$ be the 0-bridge. Clearly $\beta\subset^{\varepsilon}S$ for all $\varepsilon$ and $\beta\subset S^{*}$ for all smoothings $S^{*}$ of $S$, so by Remark 6.13 we are done. Since $T$ is of type $(0,l_{1},0,l_{2})$, the edge $H^{+}$ is outside of the tangle, i.e. $H^{+}\cap S=\varnothing$. Let AA be the connected component of $\square\setminus S$ which contains $H^{+}$, and let $C=\bar{A}\cap S$. For later use, note that it follows that
@@
We have proved Claim 6.17.1 in the case where $C$ contains no crossings of $S$. Suppose we have proved this for the case when $C$ contains $q$ crossings of $S$, and suppose that $C$ contains $q+1$ crossings. Fix a vertical splitting curve $J_{0}$ witnessing $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$. Let $x\in C\cap\operatorname{Cross}(T)$ and let $U$ be a neighborhood of $x$ witnessing the transversality (see Definition 2.1). Select $U$ to be also so small that $U\cap|J_{0}|=\varnothing$. This is possible, because by the assumption of general position, $J_{0}$ does not contain $x$ and $J_{0}$ is compact, so $d(x,J_{0})>0$. Further we can assume that $U\cap\partial\square=\varnothing$. Let $\varepsilon_{0}$ be so small that for all $\varepsilon<\varepsilon_{0}$ we can choose $U=B(x,\varepsilon)$ and also smaller than any $\varepsilon_{0}$ occurred in the previous induction steps. By the assumption that $U$ witnesses transversality, there is a homeomorphism $h\colon U\to[-1,1]^{2}$ such that $h[C]\subset([-1,1]\times\{0\})\cup(\{0\}\times[-1,1])$. Consider $h[A\cap U]$ where AA is as defined in the beginning of the proof of the lemma. By (6.32) there is a point $y\in h[A\cap U]$ such that $y$ is not in any of the coordinate axes. Without loss of generality we may assume that $y$ is in the positive-positive quarter, i.e. $y\in\left]0,1\right[^{2}$. Then we have
@@
Let $S^{\prime}$ be obtained from $S$ by replacing $S\cap U$ with $h^{-1}[\alpha^{+}\cup\alpha^{-}]$ where $\alpha^{+}=[(0,1),(1,0)]$ and $\alpha^{-}=[(-1,0),(0,-1)]$, see Figure 6. Note that one has to reparametrize the immersion of $S^{\prime}$ in an obvious way to make it a t-diagram (see Section 4.5). We did not remove any endpoints in the process, so we have $S^{\prime}\subset^{\varepsilon}_{\max}S$. Let $A^{\prime}$ and $C^{\prime}$ be the sets defined for $S^{\prime}$ in the same way as AA and $C$ were defined for $S$. It remains to show that the induction hypothesis (for $q$) can be applied to $S^{\prime}$ and $C^{\prime}$. Once we do that, for any $\varepsilon<\varepsilon_{0}$ we will obtain a maximal $\varepsilon$-smoothing of $S^{\prime}$ which contains a $(p+1)$-bridge. Since a maximal $\varepsilon$-smoothing of $S^{\prime}$ is a maximal smoothing of $S$ (Lemma 6.14), it will conclude the proof. Thus, we need to show that $\operatorname{Cross}(S^{\prime})\cap C^{\prime}$ has a smaller cardinality than $\operatorname{Cross}(S)\cap C$ and that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})=p+1$. First let us prove some preliminary claims.
@@
We will show that $\left]-1,0\right[\times\left]0,1\right[\cap hA=\varnothing$ and the other part follows by symmetry. Suppose for a contradiction that there is $z\in\left]-1,0\right[\times\left]0,1\right[\cap hA$ and let $z^{\prime}=h^{-1}(z)$. Let $y$ be as above, and $y^{\prime}=h^{-1}(y)$. Let $\gamma_{z}$ and $\gamma_{y}$ be paths connecting $z^{\prime}$ and $y^{\prime}$ respectively to $H^{+}$ and let $\varphi$ be the line segment connecting the endpoints of $\gamma_{z}$ and $\gamma_{y}$ in $H^{+}$. Let $\psi$ be the line segment connecting $z$ and $y$. If $\gamma_{z}$ and $\gamma_{y}$ intersect, modify them by perform smoothing operations within AA on their crossings to make them disjoint. Then $h^{-1}[\psi]\cup\gamma_{1}\cup\varphi\cup\gamma_{2}$ is a Jordan curve and it is entirely in $\bar{A}$. Also, the part $\gamma_{1}\cup\varphi\cup\gamma_{2}$ does not intersect $S$. But the remaining part, $h^{-1}[\psi]$ intersects $S$ only in one point. The strand of $S$ which enters the Jordan region must end in $\partial\square$, so this is a contradiction with the Jordan Curve Theorem. ∎

### v2

@@ 1 Introduction
Given a knot diagram $K$, denote by $\chi(K)$ the number of crossings in that diagram. Denote by $\mathcal{R}$ the set of classical Reidemeister moves and Reidemeister equivalence between diagrams as $K\sim_{\mathcal{R}}K^{\prime}$. Then $c(K)$ is defined as the minimim of $\{\chi(K^{\prime})\,:\,K^{\prime}\sim_{\mathcal{R}}K\}$. Now the question of additivity can be reformulated as follows. Is it possible that $K^{\prime}\sim_{\mathcal{R}}K_{0}\mathbin{\#}K_{1}$ and $\chi(K^{\prime})<c(K_{0})+c(K_{1})$? We will define a set of restricted Reidemeister moves $\mathcal{R}^{\prime}$ (in fact several versions) and show that if $K^{\prime}\sim_{\mathcal{R}^{\prime}}K$, then $\chi(K^{\prime})\geqslant c(K_{0})+c(K_{1})$.
@@ 2.1 Basics
The composition of functions is denoted $f\circ g$ but sometimes $\circ$” is ignored and it is denoted $fg$. If $A$ is a set, denote by $fA=f[A]=f(A)=|f\!\restriction\!A|=\{f(x)\,:\,x\in A\}$ the image of $A$ (usually $f[A]$ is preferred). By $\operatorname{id}_{A}$ we denote the identity map $A\to A$. The identity map fixes $A$ pointwise. Sometimes we say that a function $f$ fixes a set $A$ pointwise to mean that $f\!\restriction\!A=\operatorname{id}_{A}$.
By $\overset{\circ}{U}{}$, $\partial U$, and $\bar{U}$ denote the interior, closure, and the boundary of a set $U\subset\mathbb{R}^{n}$ respectively. By $\operatorname{Card}(A)$ denote the cardinality of the set $A$. If $\gamma_{1}$ and $\gamma_{2}$ are mappings from $I$ into some space $X$ such that $\gamma_{1}(1)=\gamma_{2}(0)$, we denote by $\gamma_{1}{}^{\frown}\gamma_{2}$ the map $\gamma\colon I\to\mathbb{R}^{2}$ given by $\gamma(t)=\gamma_{1}(2t)$ if $t<1/2$, and $\gamma(t)=\gamma_{2}(2t-1)$ otherwise. By $B(x,r)$ and $\overset{\circ}{B}{}(x,r)$ denote the closed and open balls respectively centered at $x$ with radius $r$.
@@ 2.2 Genericity
A polyhedron in $\mathbb{R}^{n}$ is a set which is the realization of a finite simplicial complex in $\mathbb{R}^{n}$ compatible with the linear structure of $\mathbb{R}^{n}$. A map $f\colon X\to Y$ between polyhedra $X$ and $Y$ is an immersion if for all $x\in X$ there is a polyhedral neighborhood $U$ of $x$ (i.e. $U$ is a polyhedron and $x\in\overset{\circ}{U}{}$) such that $f\!\restriction\!U$ is an embedding. A point $x\in|f|$ is a multiplicity point if $\operatorname{Card}(f^{-1}(x))>1$. It is a double point if $\operatorname{Card}(f^{-1}(x))=2$. Suppose $T\colon D\to\mathbb{R}^{2}$ is an immersion where $D$ is a $1$-manifold with boundary. Suppose that $x\in\mathbb{R}^{2}$ is a double point, Then $x$ is transversal if there is a compact polyhedral neighborhood $U$ of $x$, a homeomorphism $h\colon U\to[-1,1]^{2}$, and disjoint closed connected sets $C_{1},C_{2}\subset D$ each of which contains one of the points of $T^{-1}(x)$ in their interior such that $|h\circ(T\!\restriction\!C_{1})|=[-1,1]\times\{0\}$ and $|h\circ(T\!\restriction\!C_{2})|=\{0\}\times[-1,1]$.
Definition 2.2.
Let $D$ a $1$-manifold with boundary. We say that an immersion $g$ of $D$ into $\mathbb{R}^{2}$ is generic if there are at most finitely many multiplicity points, all of them are transversal double points and none of them is at $g[\partial D]$. These double points are called crossings and the set of crossings is denoted
@@ 3.3 Composition of Links
A version of Theorem 3.1 also holds for links in the following way. A link diagram is a triple $L=(L,M,\ell)$ where $L\colon D\to\square$ is a generic immersion, $M\in\mathbb{N}$, $D=I\cup(M\times S^{1})$ is the disjoint union of $I$ with $M$ many circles, $L\!\restriction\!I=(L\!\restriction\!I,\ell\!\restriction\!\operatorname{Cross}(L\!\restriction\!I))$ is a knot diagram, and for all $t\in M\times S^{1}$, $L(t)\in\overset{\circ}{\square}{}$. It is centered, if $L\!\restriction\!I$ is centered. It is composite, if $L\cap H=\{o\}$. Note that it follows that $(L\!\restriction\!I)\cap H=\{o\}$ is composite (recall the definition of $H$ in the beginning of Section 2.6). A partitioned link is a tuple $(L,M^{+},M^{-},\ell)$ such that $(L,M^{+}\cup M^{-},\ell)$ is a link and $M=M^{+}\sqcup M^{-}$. For a composite $L$, there is a unique partition $M=M^{+}\sqcup M^{-}$ such that
@@ 4.4 Applying a Homeomorphism to a Tangle
Suppose $h\colon\square\to\square$ is a homeomorphism and $(T,N,M,\ell)$ a t-diagram. Then we denote by $hT$ the t-diagram $(h\circ T,N,M,\ell\circ h^{-1})$. Important special cases are the applications of the moves $R_{0}$ and $R^{*}_{0}$, (Section 2.5).
@@ 4.5 Reparametrization
Sometimes t-diagrams need to be reparametrized. An example of this was the restriction of a t-diagram to a rectangle $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$, Definition 4.4. Another example is smoothing. A local smoothing of a t-diagram is a local change at a crossing as depicted in Figure 6. If t-diagrams $(T,N,M,\ell)$ and $(T^{\prime},N^{\prime},M^{\prime},\ell^{\prime})$ are the same t-diagram before and after smoothing, the number of strands and loops may or may not change, i.e. we might have $N\neq N^{\prime}$ and/or $M\neq M^{\prime}$, or they might not change etc. We will not give the technical details of the reparametrization and assume that the reader understands how it can be done in each specific case. A smoothing operation may connect two strands which were originally oriented in opposite directions. Then one of them changes orientation after merging with the other strand. This is not a problem, as in this context we do not care about the orientation of strands and loops in t-diagrams.
Similarly for curves. We introduced the notations $\gamma_{1}{}^{\frown}\gamma_{2}$ and $f\cup g$ in the beginning of this section, but sometimes the domain of the curves may not be exactly equal to $I$, or they may be oriented in the “wrong” direction, i.e. $\gamma_{1}(0)=\gamma_{2}(0)$ and we want to connect them at their beginning points. It should always be clear from the context how to (re-)parametrize the curves in a proper manner.
@@ 5.1 The Minimal Ray Intersection Number
A sequence of $R_{0}$-moves is equivalent to one $R_{0}$-move, because the composition of witnessing homeomorphisms is a witnessing homeomorphism. So suppose that $\widehat{K}$ has been obtained from $K$ using such a homeomorphism $h$. Recall that by the definition of the $R_{0}$-move, $h$ keeps $o$ fixed, $J$ is a ray if and only if $hJ$ is a ray and $\operatorname{\mathbf{i}}(K^{\pm},J)=\operatorname{\mathbf{i}}(hK^{\pm},hJ)$. So the statement follows from the obvious equality $hK^{\pm}=\widehat{K}^{\pm}$. ∎
@@ 6.1 Multigraphs, Matrices and Zigzag Paths
Denote by $\mathbb{N}$ natural numbers including zero and by $\mathbb{N}_{+}=\mathbb{N}\setminus\{0\}$ natural numbers excluding zero. For $n,m\in\mathbb{N}$ denote by $\llbracket n,m\rrbracket$ the discrete set $\{n,n+1,\dots,m-1,m\}$. By convention, if $m<n$, then $\llbracket n,m\rrbracket$ is the empty set and if $m=n$, then it is the singleton $\{n\}=\{m\}$. Recall that $\operatorname{Card}(A)$ the cardinality of the set $A$, e.g. for $n\leqslant m$ we have
@@
The $ij$-th element of an $(n\times n)$-matrix $e$ is expressed as $e(i,j)$ where $i,j\in\llbracket 1,n\rrbracket$. Let $\mathcal{M}^{s}_{n}(\mathbb{N})$ be the set of symmetric $\mathbb{N}$-valued $(n\times n)$-matrices,
@@
In the multigraph interpretation, $s_{i}(e)$ is the total number of out-going connections from $i$.
@@
Clearly $x_{q}>x_{q+1}$. By (6.5), ($enumi^{m}_{3}$), and the symmetry of $x_{q}$ we have $x_{q+1}\in\mathcal{M}^{s}_{n}(\mathbb{N})$. It remains to verify that $x_{q+1}$ satisfies $(*^{m^{\prime}})$. The condition $(*^{m^{\prime}}_{3})$ for $x_{q+1}$ follows trivially from the assumption that ($enumi^{m}_{3}$) holds for $x_{q}$. So let us prove $(*^{m^{\prime}}_{1})$ and $(*^{m^{\prime}}_{2})$ for $x_{q+1}$. Suppose first that $f(m)\neq m^{\prime}$. Then the definition of $x_{q+1}$ means that the number $1$ has been subtracted from exactly two rows in $x_{q}$ to obtain $x_{q+1}$. These rows are $f(m)$ and $m^{\prime}$. This means that
@@
We have proved $(*^{m^{\prime}})$ for $x_{q+1}$ in the case $m^{\prime}\neq f(m)$. Let us now assume that $f(m)=m^{\prime}$. Then the definition of $x_{q+1}$ means that the number $2$ has been subtracted frmo exactly one row of $x_{q}$ to obtain $x_{q+1}$ that row being $m^{\prime}=f(m)$. This means:
@@
Given two matrices $a$ and $b$, denote by $ab$ their Hadamard product, i.e. the matrix such that $(ab)(i,j)=a(i,j)b(i,j)$. Let us also introduce temporarily the following shorthand notation for a matrix $a$ and a set $A$: ${(A)a}=\sum_{(i,j)\in A}a(i,j)$. Then the statement (6.24) becomes
@@ 6.2 Minimal Splitting Curves
Recall Definition 4.6 of $L_{1}(T),\dots,L_{m}(T)$. Given a t-diagram $T$ and a splitting curve $J$, if the starting point of $J$ is in some $L_{i}(T)$ and the ending point is in some $L_{j}(T)$, then we call $J$ an $ij$-splitting curve (of $T$).
Definition 6.7.
Given a t-diagram $T=(T,N,M,\ell)$ denote by $\operatorname{MIN}_{ij}(T)$ the minimum number of intersections that an $ij$-splitting curve can have with $T$, formally
@@
If $m=0$ then let $\beta=\varnothing$ be the 0-bridge. Clearly $\beta\subset^{\varepsilon}S$ for all $\varepsilon$ and $\beta\subset S^{*}$ for all smoothings $S^{*}$ of $S$, so by Remark 6.13 we are done. Since $T$ is of type $(0,l_{1},0,l_{2})$, the edge $H^{+}$ is outside of the tangle, i.e. $H^{+}\cap S=\varnothing$. Let $A$ be the connected component of $\square\setminus S$ which contains $H^{+}$, and let $C=\bar{A}\cap S$. For later use, note that it follows that
@@
We have proved Claim 6.17.1 in the case where $C$ contains no crossings of $S$. Suppose we have proved this for the case when $C$ contains $q$ crossings of $S$, and suppose that $C$ contains $q+1$ crossings. Fix a vertical splitting curve $J_{0}$ witnessing $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$. Let $x\in C\cap\operatorname{Cross}(T)$ and let $U$ be a neighborhood of $x$ witnessing the transversality (see Definition 2.1). Select $U$ to be also so small that $U\cap|J_{0}|=\varnothing$. This is possible, because by the assumption of general position, $J_{0}$ does not contain $x$ and $J_{0}$ is compact, so $d(x,J_{0})>0$. Further we can assume that $U\cap\partial\square=\varnothing$. Let $\varepsilon_{0}$ be so small that for all $\varepsilon<\varepsilon_{0}$ we can choose $U=B(x,\varepsilon)$ and also smaller than any $\varepsilon_{0}$ occurred in the previous induction steps. By the assumption that $U$ witnesses transversality, there is a homeomorphism $h\colon U\to[-1,1]^{2}$ such that $h[C]\subset([-1,1]\times\{0\})\cup(\{0\}\times[-1,1])$. Consider $h[A\cap U]$ where $A$ is as defined in the beginning of the proof of the lemma. By (6.32) there is a point $y\in h[A\cap U]$ such that $y$ is not in any of the coordinate axes. Without loss of generality we may assume that $y$ is in the positive-positive quarter, i.e. $y\in\left]0,1\right[^{2}$. Then we have
@@
Let $S^{\prime}$ be obtained from $S$ by replacing $S\cap U$ with $h^{-1}[\alpha^{+}\cup\alpha^{-}]$ where $\alpha^{+}=[(0,1),(1,0)]$ and $\alpha^{-}=[(-1,0),(0,-1)]$, see Figure 6. Note that one has to reparametrize the immersion of $S^{\prime}$ in an obvious way to make it a t-diagram (see Section 4.5). We did not remove any endpoints in the process, so we have $S^{\prime}\subset^{\varepsilon}_{\max}S$. Let $A^{\prime}$ and $C^{\prime}$ be the sets defined for $S^{\prime}$ in the same way as $A$ and $C$ were defined for $S$. It remains to show that the induction hypothesis (for $q$) can be applied to $S^{\prime}$ and $C^{\prime}$. Once we do that, for any $\varepsilon<\varepsilon_{0}$ we will obtain a maximal $\varepsilon$-smoothing of $S^{\prime}$ which contains a $(p+1)$-bridge. Since a maximal $\varepsilon$-smoothing of $S^{\prime}$ is a maximal smoothing of $S$ (Lemma 6.14), it will conclude the proof. Thus, we need to show that $\operatorname{Cross}(S^{\prime})\cap C^{\prime}$ has a smaller cardinality than $\operatorname{Cross}(S)\cap C$ and that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})=p+1$. First let us prove some preliminary claims.
@@
We will show that $\left]-1,0\right[\times\left]0,1\right[\cap hA=\varnothing$ and the other part follows by symmetry. Suppose for a contradiction that there is $z\in\left]-1,0\right[\times\left]0,1\right[\cap hA$ and let $z^{\prime}=h^{-1}(z)$. Let $y$ be as above, and $y^{\prime}=h^{-1}(y)$. Let $\gamma_{z}$ and $\gamma_{y}$ be paths connecting $z^{\prime}$ and $y^{\prime}$ respectively to $H^{+}$ and let $\varphi$ be the line segment connecting the endpoints of $\gamma_{z}$ and $\gamma_{y}$ in $H^{+}$. Let $\psi$ be the line segment connecting $z$ and $y$. If $\gamma_{z}$ and $\gamma_{y}$ intersect, modify them by perform smoothing operations within $A$ on their crossings to make them disjoint. Then $h^{-1}[\psi]\cup\gamma_{1}\cup\varphi\cup\gamma_{2}$ is a Jordan curve and it is entirely in $\bar{A}$. Also, the part $\gamma_{1}\cup\varphi\cup\gamma_{2}$ does not intersect $S$. But the remaining part, $h^{-1}[\psi]$ intersects $S$ only in one point. The strand of $S$ which enters the Jordan region must end in $\partial\square$, so this is a contradiction with the Jordan Curve Theorem. ∎